OpenStudy (hnamowicz):

Which is an equation in point-slope form for the line that passes through the points (−1,4)and (3,−4) ?

OpenStudy (hnamowicz):

y+4=2(x−3)y+4=2(x−3) y+4=−2(x−3)y+4=−2(x−3) y−3=−2(x+4)y−3=−2(x+4) y−4=2(x+3)

OpenStudy (3mar):

Could I help you? Are you familiar with that?